# Leveraging Large Language Models (LLMs) to Empower Training-Free Dataset Condensation for Content-Based Recommendation

### v3

@@ IV-E Further Investigation
In this subsection, we further investigate how different values of hyperparameter $\alpha$ affect the quality of the condensed datasets. The hyperparameter $\alpha$ is utilized to balance the influence of user interests and user embeddings on the user synthesis. The larger value of $\alpha$ denotes larger influence of user interests on the synthesis of fake users. As shown in Figure 5, we can find that the best choice for $\alpha$ with three different recommendation models is smaller than 1 but larger than 0, which verifies that the effectiveness of selecting/merging users with similar interests for the synthesis.
V Related Work
V-A Content-Based Recommendation
Content-based recommendation (CBR) involves a wide range of information in different formats: text, image, video, social networks, etc [32]. In this paper, we mainly focus on textual contents, i.e., news, books, and movie reviews.
In recent years, textual CBR has garnered substantial attention and has been the subject of comprehensive investigation within both industrial and academic domains. Models based on neural networks [4, 5] have been proposed to learn the representations of users and items via CNN [33] or attention mechanism [34]. Thereafter, the advance of pretrained language models like BERT [35] exhibited promising performance on textual content recommendation [36, 37]. As those methodologies depend on training with extensive-scale datasets, their exceptional performance comes at a high computational cost. The recent emergence of inquiries into dataset condensation and dataset distillation presents a promising avenue for tackling this concern [7, 11]. The inherent concept underlying dataset condensation and dataset distillation revolves around the synthesis of a compact dataset for model training, with the objective of enabling models trained on this condensed dataset to achieve performance levels comparable to those trained on the complete dataset. Nevertheless, no extant research has delved into the realm of condensation specifically within the context of CBR. To bridge this critical gap, we introduce an innovative and efficacious framework tailored to the task of dataset condensation for CBR.
V-B Dataset Condensation
It is notorious that training neural networks on large datasets is prohibitively expensive. To mitigate this issue, plenty of works on dataset distillation/dataset condensation have been proposed to synthesize small datasets to replace the large ones, such as GCond [8], DC [11] and SFGC [9]. Those methods always formulate the condensation process as a bi-level optimization problem. The condensation methods have been explored on images [11], graph [8] and text embeddings [12]. Various methods have been proposed to improve the quality of condensation and optimization efficiency. For instance, DC [11] proposed gradient matching, achieving a comparable performance under a high compression ratio, and DosCond [10] proposed one-step update strategy, which significantly improves the optimization efficiency. Condensation on textual data is conducted based on the text embeddings or soft labels [38, 39, 40, 13, 12]. Further exploration extends condensation methods to image-text retrival [41]. Recently, DConRec [15] is proposed to condense the datasets for ID-based recommendation and CGM [16] is proposed to handle the datasets for CTR prediction with side information. Distill-CF [14] proposes a condensation method tailored for NTK-based models.
However, rare attention has been paid to the domain of CBR. To address this disparity, we posit the consolidation of the CBR dataset through the utilization of Large Language Models (LLM). Unlike earlier methodologies, our dataset condensation approach is devoid of training requirements, leading to a substantial reduction in condensation expenses. Furthermore, our condensation procedure centers on textual content, thereby diverging from those text condensation methods that hinge on text embeddings.
VI Conclusion
@@
Advanced content-based recommendation models rely on training with large datasets, which is highly expensive. To solve this problem, in this paper, we investigate how to condense the dataset of content-based recommendation into a small synthetic dataset, trained on which models can achieve comparable performance with those trained on the full dataset. Consequently, we propose a novel method TF-DCon that utilizes LLMs to condense contents of items and users, where users and their corresponding historical interactions are condensed via a curated clustering-based synthesis module. To the best of our knowledge, this is the first work on condensing datasets for content-based recommendation and the first condensation method that is free from iterative optimizing synthetic data. The experimental results on multiple real-world datasets verify the effectiveness of our proposed methods. Notably, we approximate up to $97\%$ of the original performance while reducing the dataset size by $95\%$ (i.e., on dataset MIND). The model training on the condensed datasets is significantly faster (i.e., 5×5$\times$5 × speedup).

### v4

@@ 5.7 Further Investigation
In this section, we further investigate how different values of hyperparameter $\alpha$ affect the quality of the condensed datasets. The hyperparameter $\alpha$ is utilized to balance the influence of user interests and user embeddings on the user synthesis. The larger value of $\alpha$ denotes larger influence of user interests on the synthesis of fake users. As shown in Figure 5, we can find that in most cases, the best choice for $\alpha$ with three different recommendation models is smaller than 1 but larger than 0, which verifies that the effectiveness of selecting/merging users with similar interests for the synthesis. Besides, the selection of $\alpha$ varies among three recommendation models, influenced by their respective interest-capturing ability. A smaller optimal $\alpha$ indicates the model’s diminished capacity to capture information from limited texts, such as user interests. We can observe that Fastformer possesses the best interest-capturing ability, which is also in line with the results in previous study Wu et al. (2021).
5.7.2 Analysis on the value of K
In this section, we study how the value of $K$ affect the condensation performance. To be noted, the $K$ equals the number of synthesized users in the condensed dataset. Therefore, increasing the number of $K$ will increase the size of the condensed dataset. Specifically, we conduct experiment on three datasets with the recommendation model NAML. The result is reported in Table 8, where we can observe that increasing $K$ can lead to better performance of NAML, denoting better quality of the condensed dataset.
5.7.3 Effectiveness of User Interests
[…]
Table 11: Ablation Study on User Interests.
Dataset Rec Model Metric N@1 N@5 R@1 R@5
[…]
TF-DCon 0.3071 0.3691 0.4377 0.6150
[…]
TF-DCon 0.2997 0.3597 0.4279 0.6000
[…]
TF-DCon 0.3022 0.3637 0.4334 0.6096
[…]
TF-DCon 0.5411 0.8033 0.4704 0.9984
[…]
TF-DCon 0.5453 0.8054 0.4751 0.9985
[…]
TF-DCon 0.5548 0.8092 0.4851 0.9986
[…]
TF-DCon 0.8484 0.8494 0.1873 0.7475
[…]
TF-DCon 0.8149 0.8371 0.1798 0.7446
[…]
TF-DCon 0.8251 0.8429 0.1836 0.7465
5.7.4 Moderation Check.
To check whether the condensed dataset contain toxic contents, we conduct moderation check via the endpoints provided by OpenAI [^1] https://platform.openai.com/docs/guides/moderation. Specifically, we randomly sample 100 items from each dataset to assess their contents. The results are presented in table 12, where the "Mean" signifies the mean violation score of the contents in items encompassing the corresponding toxic words and "Var" denotes the associated variance. The violation score represents the model’s confidence in input violations of the OpenAI category policy. This value ranges between 0 and 1, with higher values indicating higher confidence that the input contains the corresponding toxic words. We can observe from the table that the condensed datasets derived from the three original datasets do not incorporate any of the toxic contents.
Table 12: Moderation Check.
Dataset MIND Goodreads MovieLens
Violation Category Mean Var Mean Var Mean Var
sexual 4.94E-03 2.13E-03 8.45E-04 1.43E-05 3.08E-04 1.13E-06
hate 6.72E-05 3.97E-08 3.33E-05 1.90E-08 2.11E-05 3.75E-09
harassment 1.57E-04 2.64E-07 4.55E-05 6.64E-09 4.64E-05 4.29E-08
self-harm 2.46E-06 8.65E-11 3.04E-06 1.77E-10 2.71E-06 8.68E-11
sexual/minors 1.05E-05 4.26E-09 1.55E-05 1.09E-08 2.68E-06 5.28E-11
hate/threatening 1.18E-06 2.12E-11 1.01E-06 1.70E-11 1.11E-06 2.10E-11
violence/graphic 4.77E-04 4.31E-06 7.49E-05 9.46E-08 5.03E-05 3.11E-08
self-harm/intent 2.43E-07 1.43E-12 7.80E-08 1.18E-13 1.38E-07 3.61E-13
self-harm/instructions 7.96E-07 6.16E-12 3.07E-08 3.02E-14 9.53E-08 2.98E-13
harassment/threatening 2.13E-05 1.36E-08 1.75E-06 2.60E-11 4.18E-06 4.96E-10
violence 3.40E-03 2.57E-04 4.79E-03 3.50E-04 3.42E-03 1.10E-04
5.8 Case Study
In this section, we conduct case study to examine the reasonability of the condensed data by investing the correlation between the history of the synthesized users and their corresponding real users. The following example is randomly sampled from the MovieLens dataset. We show the history of the synthesized user and three related real users. In the following, the history of synthesized user consists of short introduction of the interacted movies while the histories of real users consist of the title and description of the interacted movies. We can observe that the preference information of those real users can be roughly covered by this synthesized user, which may be the reason why models trained on the condensed data since it provides adequate information for sufficient model training.
Synthesized User Interaction History: My Best Friend’s Wedding is about a romantic comedy.  from: //Real User 2 Back to the Future is about a high school student who time travels to the past and must ensure his parents fall in love to preserve his own existence.  from: //Real User 2 Grosse Pointe Blank is about a hitman facing unexpected dilemmas at a high school reunion.  from: //Real User 3 "Robin Hood: Men in Tights" is about a skilled archer leading a funny rebellion against a corrupt ruler, with the help of his misfit Merry Men.  from: //Real User 1 & Real User 2 & Real User 3 Mad City is about a hostage situation in a museum, delving into the power struggle between a security guard and a journalist, while examining media ethics and the blurred boundaries of truth.  from: //Real User 1 The Wizard of Oz is about Dorothy’s journey in the magical land of Oz, where she seeks the help of the Great Wizard to return home, facing challenges and learning valuable lessons along the way.  from: //Real User 1 This Is Spinal Tap is a mockumentary comedy released in 1984.  from: //Real User 1
Real User 1 Interaction History: Snow White and the Seven Dwarfs (1937), with description: "Snow White and the Seven Dwarfs is a timeless animated film released in 1937. It tells the story of a young princess named Snow White who finds herself in the care of seven lovable dwarfs after escaping from her evil stepmother. With its enchanting tale, memorable songs, and captivating animation, this classic movie continues to captivate audiences of all ages." Robin Hood: Men in Tights (1993), with description: "Robin Hood: Men in Tights (1993), a comedic retelling of the classic Robin Hood tale, a skilled archer named Robin must outwit the corrupt Prince John and win the heart of the beautiful Maid Marian. With a band of misfit Merry Men by his side, Robin leads a hilarious and daring rebellion against the unjust ruler, using wit, skill, and plenty of laughs along the way." The Wizard of Oz (1939), with description: "The Wizard of Oz is a classic fantasy film released in 1939. It tells the story of a young girl named Dorothy, who is transported to the magical land of Oz after a tornado hits her Kansas farm. In Oz, Dorothy encounters various colorful characters, including the Scarecrow, the Tin Man, and the Cowardly Lion, as she embarks on a journey to find the Great Wizard who can help her return home. Along the way, they face challenges from the Wicked Witch of the West, and learn important lessons about friendship, courage, and the importance of following your dreams. The Wizard of Oz is beloved for its stunning visuals, memorable songs, and timeless themes." Mad City (1997), with description: "Mad City, a 1997 film, centers on a hostage crisis in a museum, delving into the power struggle between a disgruntled security guard and a manipulative journalist, each vying to control the crisis narrative. Through intense dialogue and suspense, Mad City critically examines media ethics and the blurred boundaries between truth and fiction." This Is Spinal Tap (1984), with description: "This Is Spinal Tap is a comedy mockumentary film released in 1984."
Real User 2 Interaction History: My Best Friend’s Wedding (1997), with description: "My Best Friend’s Wedding is a romantic comedy film released in 1997." Snow White and the Seven Dwarfs (1937), with description: "Snow White and the Seven Dwarfs is a timeless animated film released in 1937. It tells the story of a young princess named Snow White who finds herself in the care of seven lovable dwarfs after escaping from her evil stepmother. With its enchanting tale, memorable songs, and captivating animation, this classic movie continues to captivate audiences of all ages." Die Hard (1988), with description: "Die Hard is an action film released in 1988." Robin Hood: Men in Tights (1993), with description: "In Robin Hood: Men in Tights (1993), a comedic retelling of the classic Robin Hood tale, a skilled archer named Robin must outwit the corrupt Prince John and win the heart of the beautiful Maid Marian. With a band of misfit Merry Men by his side, Robin leads a hilarious and daring rebellion against the unjust ruler, using wit, skill, and plenty of laughs along the way." Back to the Future (1985), with description: "Back to the Future is a science fiction film released in 1985. The movie follows the story of a high school student named Marty McFly who, with the help of a eccentric scientist, travels back in time to the year 1955. Marty must find a way to return to the present and also ensure that his parents meet and fall in love, or else he risks erasing his own existence. The film combines comedy, adventure, and a touch of romance as it explores the consequences of altering the past and the importance of family."
Real User 3 Interaction History: Robin Hood: Men in Tights (1993), with description: "Robin Hood: Men in Tights" (1993), a comedic retelling of the classic Robin Hood tale, a skilled archer named Robin must outwit the corrupt Prince John and win the heart of the beautiful Maid Marian. With a band of misfit Merry Men by his side, Robin leads a hilarious and daring rebellion against the unjust ruler, using wit, skill, and plenty of laughs along the way." Chasing Amy (1997), with description: "‘Chasing Amy’ is a 1997 romantic comedy-drama film." In the Name of the Father (1993), with description: "In the Name of the Father is a gripping drama film released in 1993. It tells the story of Gerry Conlon, a young Irishman wrongfully convicted of an IRA bombing. Set in the backdrop of the Troubles in Northern Ireland, the film explores themes of injustice, incarceration, and the quest for truth and justice. Directed by Jim Sheridan and featuring remarkable performances by Daniel Day-Lewis and Pete Postlethwaite, this emotional and thought-provoking movie delves into the complexities of the legal system and the enduring power of family bonds." Nightmare on Elm Street, A (1984), with description: "A Nightmare on Elm Street is a horror film released in 1984. Directed by Wes Craven, the movie tells the story of a group of teenagers who are haunted in their dreams by a supernatural killer named Freddy Krueger. As they struggle to stay awake and escape his clutches, they soon realize that death in their dreams also means death in reality. Filled with suspenseful moments and unnerving visual effects, A Nightmare on Elm Street delivers a chilling and unforgettable experience for horror enthusiasts." Grosse Pointe Blank (1997), with description: "Grosse Pointe Blank is a 1997 movie that revolves around a skilled and professional hitman who, during a high school reunion, unexpectedly finds himself confronting personal and moral dilemmas."
5.9 Implications
The results show that that our proposed method is effective in condense content-based recommendation datasets, achieving comparable model training performance with even $5\%$ data. Our method designs a novel condensing paradigm, training-free, from previous iterative-optimization based ones. Benefiting from this design, our method is significantly more efficient. Besides, we enable the generation of discrete texts via ChatGPT and preserve users’ preference via user-level condensation module. Therefore, incorporating our method to condense datasets for daily model training can save lots of resources costs without performance degradation in practice.
6 Limitations
Although TF-DCon has been introduced to address the issue of resources demanding during model training by condensing datasets, it introduces new limitations: (1) Although the condensation process is efficient, it requires a pre-process of the dataset by ChatGPT/other LLMs, the efficiency of which can be unstable when we utilize ChatGPT for the pre-process since it is affected by the network latency. (2) Since the cost of calling ChatGPT is linear to the size of dataset, we may need to train the models based on the condensed datasets for a certain times to achieve the cost-gain balance. (3) As shown in the experiments, the performance of TF-DCon with open source LLMs is still limited.
7 Conclusion
@@
Advanced content-based recommendation models rely on training with large datasets, which is costly. To solve this problem, in this paper, we investigate how to condense the dataset of content-based recommendation into a small synthetic dataset, enabling models to achieve performance comparable to those trained on full datasets. We propose a novel method TF-DCon where a prompt-evolving module is proposed to adapt ChatGPT to condense contents of items, and users and their corresponding historical interactions are condensed via a curated clustering-based synthesis module. This work represents the first exploration of dataset condensation for content-based recommendation and the first non-iterative synthetic data optimization approach. The experimental results on multiple real-world datasets verify the effectiveness of our proposed methods. Notably, we approximate up to $97\%$ of the original performance while reducing the dataset size by $95\%$ (i.e., on the MIND dataset). The condensation process of TF-DCon is significantly faster than the method that is based on nested optimization.
CRediT Authorship Contribution Statement
Jiahao Wu: Conceptualization, Methodology, Partial experiments, Validation, Formal analysis, Investigation, Manuscript writing. Qijiong Liu: Data Curation, Main experiments, Validation, Formal analysis, Investigation. Hengchang Hu: Formal Analysis, Review, Editing. Wenqi Fan: Conceptualization, Formal Analysis, Review, Supervision. Shengcai Liu: Formal Analysis, Review, Editing, Supervision. Qing Li: Conceptualization, Formal Analysis, Review, Supervision. Xiaoming Wu: Conceptualization, Formal Analysis, Review, Supervision. Ke Tang: Conceptualization, Formal Analysis, Review, Editing, Supervision.
Resource Availability
Code and data will be made available on request.
Declaration of Generative AI and AI-assisted Technologies in the Writing Process
During the preparation of this work the author(s) used ChatGPT in order to to improve language and readability, with caution. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the publication.